# The geometry of preperiodic points in families of maps on $\mathbb{P}^{N}$

### v1

@@ Abstract.
We study the dynamics of algebraic families of maps on $\mathbb{P}^{N}$, over the field $\mathbb{C}$ of complex numbers, and the geometry of their preperiodic points. The goal of this note is to formulate a conjectural characterization of the subvarieties of $S\times\mathbb{P}^{N}$ containing a Zariski-dense set of preperiodic points, where the parameter space $S$ is a quasiprojective complex algebraic variety; the characterization is given in terms of the non-vanishing of a power of the invariant Green current associated to the family of maps. This conjectural characterization is inspired by and generalizes the Relative Manin-Mumford Conjecture for families of abelian varieties, recently proved by Gao and Habegger, and it includes as special cases the Manin-Mumford Conjecture (theorem of Raynaud) and the Dynamical Manin-Mumford Conjecture (posed by Ghioca, Tucker, and Zhang). We provide examples where the equivalence is known to hold, and we show that several recent results can be viewed as special cases. Finally, we give the proof of one implication in the conjectural characterization.
@@ 1. Introduction
given by $\Phi(s,z)=(s,f_{s}(z))$, where each $f_{s}$ is an endomorphism of the complex projective space $\mathbb{P}^{N}$ of degree $d$. Throughout this note, we let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ denote a closed irreducible subvariety which is flat over $S$. We will use boldface $\mathbf{X}$ to denote the generic fiber of $\mathcal{X}$ and let $\mathbf{\Phi}:\mathbf{P^{N}}\to\mathbf{P^{N}}$ be the map induced by $\Phi$, viewed as an endomorphism over the function field $\mathbb{C}(S)$.
@@
This is an ambitious goal; even the case where $S$ is a point remains vastly open, where a conjectural characterization goes by the name of the Dynamical Manin-Mumford Conjecture, proposed by Shouwu Zhang [Zh] and reformulated in [GTZ] and [GT]. Roughly speaking, if $\dim S=0$, the subvariety $\mathcal{X}$ of $\mathbb{P}^{N}$ should contain a Zariski-dense set of preperiodic points if and only if it is itself preperiodic (or is preperiodic for an endomorphism that commutes with $\Phi$). For $\dim S>0$, we will see that the variety $\mathcal{X}$ needs only be “big enough” in a preperiodic subvariety.
To make a precise conjecture, we employ the notion of $\Phi$-special subvarieties of Ghioca and Tucker in [GT]. We say that an irreducible subvariety $\mathcal{Y}\subset S\times\mathbb{P}^{N}$, which is flat over a Zariski-open subset of $S$ and Zariski-closed in $S\times\mathbb{P}^{N}$, is $\Phi$-special if there exist a subvariety $\mathbf{Z}\subset\mathbf{P^{N}}$ over the algebraic closure $\overline{\mathbb{C}(S)}$ containing the generic fiber ${\bf Y}$, a polarizable endomorphism $\mathbf{\Psi}:\mathbf{Z}\to\mathbf{Z}$, and an integer $n\in\mathbb{N}$ so that the following hold:
@@
Recall that an endomorphism $\Psi$ of a projective variety $Z$ is polarizable if there exist an ample line bundle $L$ on $Z$ and integer $q>1$ so that $\Psi^{*}L\simeq L^{q}$. Polarizable is equivalent to being the restriction of an endomorphism $\tilde{\Psi}:\mathbb{P}^{N}\to\mathbb{P}^{N}$ to an invariant subvariety, where the embedding $Z\hookrightarrow\mathbb{P}^{N}$ may be defined by a power of $L$ and $\Psi=\tilde{\Psi}|_{Z}$; see, for example, [Fa1] [MZ1] [MZ2] for more on polarized endomorphisms.
@@
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree $>1$, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be a complex, irreducible subvariety which is flat over $S$. The following are equivalent.
@@
Here $\hat{T}_{\Phi}$ is the canonical Green current associated to $\Phi$ on $S\times\mathbb{P}^{N}$. We briefly recall its construction, as given for instance in [GV, §2.3]. Let $\omega$ be the Fubini-Study form on $\mathbb{P}^{N}$, and let $\hat{\omega}$ be the smooth $(1,1)$-form on $S\times\mathbb{P}^{N}$ defined by pulling back $\omega$ via the projection $S\times\mathbb{P}^{N}\to\mathbb{P}^{N}$. Then the sequence $d^{-n}(\Phi^{n})^{*}(\hat{\omega})$ converges weakly to the closed positive $(1,1)$-current $\hat{T}_{\Phi}$ on $S\times\mathbb{P}^{N}$; the potentials converge locally uniformly. If $\mathcal{Y}\subset S\times\mathbb{P}^{N}$ is a $\Phi$-special subvariety which is invariant for an endomorphism $\Psi$ that commutes with an iterate of $\Phi$, then $\hat{T}_{\Phi}=\hat{T}_{\Psi}$ on $\mathcal{Y}$ because of the commuting relation. Moreover, the slices of $\hat{T}_{\Phi}^{\wedge\dim_{S}\mathcal{Y}}$ in fibers of $\mathcal{Y}\to S$ are the measures of maximal entropy for the restriction of $\Psi$; see, for example, [DS2]. So condition (2) means that (an iterate of) $\mathcal{X}$ intersects these families of measures nontrivially. By convention, $\hat{T}^{\wedge 0}=1$. We point out here that the non-vanishing of $\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}\wedge[\mathcal{X}]$ can be thought of from an arithmetic viewpoint in view of the theory of adelic metrized line bundles on quasi-projective varieties developed by Yuan–Zhang [YZ1]. In the notation of [Guo] and [YZ1], it is equivalent to the non-vanishing of an intersection number $\tilde{L_{\Phi}}^{\wedge r_{\Phi,\mathcal{X}}}|_{X}\cdot\overline{H}^{\dim\mathcal{X}-r_{\Phi,\mathcal{X}}}$ for some $\overline{H}\in\widetilde{\mathrm{Pic}}(X/\mathbb{C})$ as shown in [Guo, Theorem 1.2].
@@
Conjecture 1.1 is known to hold in all cases when $N=1$, as we explain in Section 3. Conjecture 1.1 is also known to hold in some settings when $\dim S=0$ and $N>1$, moving beyond the cases where $\Phi$ preserves an abelian variety; see §3.1. Moreover, in the following sections, we explain that many existing works and conjectures can be viewed as special cases of Conjecture 1.1.
A weaker form of Conjecture 1.1 can be formulated as follows:
@@
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be a family of morphisms of degree $>1$, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible, flat family of subvarieties over $S$ containing a Zariski-dense set of preperiodic points of $\Phi$. Then $\mathcal{X}$ has codimension $\leq\dim S$ in a $\Phi$-special subvariety.
@@
Gao and Habegger pointed out in [GH2] that the converse assertion to Conjecture 1.2 is not true, with explicit examples and constructions in the setting where $\Phi$ fixes an abelian subscheme of $S\times\mathbb{P}^{N}$; see also [Gao]. This led them to formulate their [GH2, Theorem 1.3], which is what our Conjecture 1.1 aims to extend. Some of those counterexample constructions arose already in the (conjectural) characterization of subvarieties with geometric canonical height 0, in the Geometric Dynamical Bogomolov Conjecture formulated by Gauthier and Vigny [GV]. The conjecture in [GV] aims to extend theorems for abelian varieties over function fields of characteristic 0, proved in [GH1, CGHX]. We discuss the relation between [GV, Conjecture 1.9] and Conjecture 1.1 in Section 5.
We complete this note with a proof of the following implication.
@@
Remark 1.4.
In the case that both $\Phi$ and $\mathcal{X}$ are defined over $\overline{\mathbb{Q}}$, we expect Conjectures 1.1 and 1.2 to remain true upon replacing $\Phi$-preperiodic points by $\Phi$-small points. Here a sequence of points $\{x_{n}\}_{n}\subset\mathcal{X}(\overline{\mathbb{Q}})$ is called $\Phi$-small if $\hat{h}_{\Phi}(x_{n})\to 0$, where $\hat{h}_{\Phi}(x):=\hat{h}_{\Phi_{\pi(x)}}(x)$ is the fiber-wise Call–Silverman canonical height introduced in [CS], and $\pi:S\times\mathbb{P}^{N}\to S$ is the projection.
Contents of this article. We begin by introducing a dynamical notion of $\Phi$-rank of $\mathcal{X}$, and we compare Conjecture 1.1 to the theorems of Gao and Habbeger from [GH2]. In particular, Conjecture 1.1 characterizes subvarieties $\mathcal{X}\subset S\times\mathbb{P}^{N}$ of maximal $\Phi$-rank; see Conjecture 2.3. We also compare the conditions of Conjecture 1.1 to the concept of non-degeneracy of subvarieties $\mathcal{X}$, which was originally introduced for families of abelian varieties. We then provide a brief survey of known cases of Conjecture 1.1, outside of the setting of abelian schemes. In particular, we remind the reader of the Dynamical Manin-Mumford Conjecture, first posed in [Zh], which corresponds to Conjecture 1.1 in the case where $\dim S=0$. We show that Conjecture 1.1 is known to hold in dimension $N=1$ over any base $S$ (and is in fact equivalent to [De3, Theorem 1.1]), and we relate it to the concept of $J$-stability for maps on $\mathbb{P}^{1}$. We then show that a conjecture of [BD2], and the recent classification of “special curves” in the moduli space $\mathrm{M}_{d}^{1}$ of maps on $\mathbb{P}^{1}$ in [JX], is a special case of Conjecture 1.1; see Theorem 3.5. Moreover, we illustrate the strength of Conjecture 1.1 with an example showing that uniform versions of the conjecture are consequences of the conjecture itself; the example we provide comes from the study of shared preperiodic points for distinct maps on $\mathbb{P}^{1}$. For maps on $\mathbb{P}^{N}$, with $N>1$, we show that Conjecture 1.1 implies the recent sparsity theorem of Gauthier, Taflin, and Vigny in [GTV] about PCF maps in the moduli space $\mathrm{M}_{d}^{N}$ of maps on $\mathbb{P}^{N}$. We explore the concept of minimal $\Phi$-rank and compare Conjecture 1.1 to the Geometric Bogomolov Conjecture posed in [GV]. Finally, in Section 6, we provide the proof of Theorem 1.3.
Acknowledgements. We would like to thank Ziyang Gao, Thomas Gauthier, Lars Kühne, Harry Schmidt, and Gabriel Vigny for many interesting discussions and their help during the preparation of this note. We also thank the Simons Foundation for their support during a Symposium in August 2022 where this work was initiated. This project was supported in part with funding from the National Science Foundation and the Radcliffe Institute for Advanced Study.
2. $\Phi$-rank and Betti-rank
@@ 2.1. $\Phi$-Rank
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree $>1$, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible flat subvariety over $S$. We define the $\Phi$-rank of $\mathcal{X}$ to be
@@
For $\dim S=0$, so that $\Phi$ is a single endomorphism $f:\mathbb{P}^{N}\to\mathbb{P}^{N}$, and for any irreducible subvariety $Z$ of $\mathbb{P}^{N}$, we have
@@
[…]
@@
$\dim_{S}\mathcal{X}\leq\operatorname{rank}_{\Phi}(\mathcal{X})\leq r_{\Phi,\mathcal{X}}$
for the relative special dimension $r_{\Phi,\mathcal{X}}$ of $\mathcal{X}$.
@@
Suppose that $\mathcal{Y}$ is $\Phi$-special, and set $k=\dim_{S}\mathcal{Y}$. Let $\Psi$ be a family of polarized endomorphisms over $S$ that commutes with an iterate of $\Phi$ along $\mathcal{Z}$ with $\mathcal{Y}\subset\mathcal{Z}\subset S\times\mathbb{P}^{N}$; replacing $\Psi$ with an iterate, we assume that $\mathcal{Y}$ is prefixed for $\Psi$. Now replace $\mathcal{Y}$ with an iterate under $\Psi$ so that $\Psi(\mathcal{Y})=\mathcal{Y}$, and pass to a normalization if it is not itself normal. We have $\hat{T}_{\Phi}=\hat{T}_{\Psi}$ on $\mathcal{Y}$. Since $(\Psi,\mathcal{Y})$ defines a family of polarized dynamical systems of degree $e>1$, there is a (1,1)-form $\omega$ on $\mathcal{Y}$ representing the first Chern class of the polarizing (relatively ample) line bundle $\mathcal{L}$ so that $\omega^{k+1}=0$. Because of the local-uniform convergence of the potentials of pullbacks of $\omega$ to a potential for $\hat{T}_{\Psi}$, we know that the wedge power $\hat{T}_{\Psi}^{\wedge j}$ is the limit of pullbacks $\frac{1}{e^{jn}}(\Psi^{n})^{*}(\omega^{\wedge j})$ on $\mathcal{Y}$, for any $j\geq 1$. As $\omega^{\wedge(k+1)}=0$, we have $\hat{T}_{\Phi}^{\wedge(1+\dim_{S}\mathcal{Y})}\wedge[\mathcal{Y}]=0$.
@@
For the lower bound on $\operatorname{rank}_{\Phi}(\mathcal{X})$, observe that the slice $(1,1)$-current $T_{s}$ of $\hat{T}_{\Phi}$ over $s\in S$ satisfies $T_{s}^{\wedge\ell}\wedge[X]\not=0$ for every subvariety $X$ of dimension $\ell$ in $\mathbb{P}^{N}$ and every $s\in S$, by Lemma 2.1. The nonvanishing of $\hat{T}_{\Phi}^{\wedge\dim_{S}\mathcal{X}}\wedge[\mathcal{X}]$ follows; see, for example, [BB, Proposition 4.3] on slicing. ∎
By Proposition 2.2, Conjecture 1.1 becomes a characterization of subvarieties $\mathcal{X}$ of maximal $\Phi$-rank, in the following sense:
@@
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree $>1$, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible subvariety which is flat over $S$. The following are equivalent:

### v2

@@ Abstract.
We study the dynamics of complex algebraic families of maps on $\mathbb{P}^{N}$, and the geometry of their preperiodic points. The goal of this article is to formulate a conjectural characterization of the subvarieties of $S\times\mathbb{P}^{N}$ containing a Zariski-dense set of preperiodic points, where the parameter space $S$ is a quasiprojective complex algebraic variety; the characterization is given in terms of the non-vanishing of a power of the invariant Green current associated to the family of maps. This conjectural characterization is inspired by and generalizes the Relative Manin-Mumford Conjecture for families of abelian varieties, recently proved by Gao and Habegger, and it includes as special cases the Manin-Mumford Conjecture (theorem of Raynaud) and the Dynamical Manin-Mumford Conjecture (posed by Ghioca, Tucker, and Zhang). We provide examples where the equivalence is known to hold, and we show that many recent results can be viewed as special cases. Finally, we give the proof of one implication in the conjectural characterization.
Key words and phrases:
preperiodic points, Manin–Mumford Conjecture, Green currents, bifurcation
1991 Mathematics Subject Classification:
37F46, 37P35
@@ 1. Introduction
given by $\Phi(s,z)=(s,f_{s}(z))$, where each $f_{s}$ is an endomorphism of the complex projective space $\mathbb{P}^{N}$ of degree $d$. Throughout this article, we let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ denote a closed irreducible subvariety which is flat over $S$. We will use boldface $\mathbf{X}$ to denote the generic fiber of $\mathcal{X}$ and let $\mathbf{\Phi}:\mathbf{P^{N}}\to\mathbf{P^{N}}$ be the map induced by $\Phi$, viewed as an endomorphism over the function field $\mathbb{C}(S)$.
@@
This is an ambitious goal; even the case where $S$ is a point remains vastly open, where a conjectural characterization goes by the name of the Dynamical Manin-Mumford Conjecture, proposed by Shouwu Zhang [Zh] and reformulated in [GTZ] and [GT]. Roughly speaking, if $\dim S=0$, a subvariety $X$ of $\mathbb{P}^{N}$ should contain a Zariski-dense set of preperiodic points if and only if it is itself preperiodic (or is preperiodic for an endomorphism that commutes with $\Phi$). For $\dim S>0$, we will see that the variety $\mathcal{X}\subset S\times\mathbb{P}^{N}$ needs only be “big enough” in a preperiodic subvariety.
To make a precise conjecture, we employ the notion of $\Phi$-special subvarieties introduced by Ghioca and Tucker in [GT]. We say that an irreducible subvariety $\mathcal{Y}\subset S\times\mathbb{P}^{N}$, which is flat over a Zariski-open subset of $S$ and Zariski-closed in $S\times\mathbb{P}^{N}$, is $\Phi$-special if there exist a subvariety $\mathbf{Z}\subset\mathbf{P^{N}}$ over the algebraic closure $\overline{\mathbb{C}(S)}$ containing the generic fiber $\bf Y$} of $\mathcal{Y}$, a polarizable endomorphism $\mathbf{\Psi}:\mathbf{Z}\to\mathbf{Z}$, and an integer $n\in\mathbb{N}$ so that the following hold:
@@
An endomorphism $\Psi$ of a projective variety $Z$ (over any field of characteristic 0) is polarizable if there exist an ample line bundle $L$ on $Z$ and integer $q>1$ so that $\Psi^{*}L\simeq L^{q}$. Polarizable is equivalent to being the restriction of an endomorphism $\tilde{\Psi}:\mathbb{P}^{N}\to\mathbb{P}^{N}$ to an invariant subvariety, where the embedding $Z\hookrightarrow\mathbb{P}^{N}$ may be defined by a power of $L$ and $\Psi=\tilde{\Psi}|_{Z}$; see, for example, [Fa1] [MZ1] [MZ2] for more on polarized endomorphisms.
@@
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree >1>1, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be a complex, irreducible subvariety which is flat over $S$. The following are equivalent.
@@
Here, by convention, $\hat{T}^{\wedge 0}=1$. Note that if $\mathcal{Y}\subset S\times\mathbb{P}^{N}$ is a $\Phi$-special subvariety which is invariant for an endomorphism $\Psi$ that commutes with an iterate of $\Phi$, then $\hat{T}_{\Phi}=\hat{T}_{\Psi}$ on $\mathcal{Y}$ because of the commuting relation. Moreover, the slices of $\hat{T}_{\Phi}^{\wedge\dim_{S}\mathcal{Y}}$ in fibers of $\mathcal{Y}\to S$ are the measures of maximal entropy for the restriction of $\Psi$; see, for example, [DS2]. So condition (2) of Conjecture 1.1 means that (an iterate of) $\mathcal{X}$ intersects these families of measures nontrivially. We point out here that the non-vanishing of $\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}\wedge[\mathcal{X}]$ can also be seen from an arithmetic viewpoint in the theory of adelic metrized line bundles on quasi-projective varieties developed by Yuan–Zhang [YZ1]. In the notation of [Guo] and [YZ1], it is equivalent to the non-vanishing of an intersection number $\widetilde{L_{\Phi}}^{r_{\Phi,\mathcal{X}}}|_{X}\cdot\overline{H}^{\dim\mathcal{X}-r_{\Phi,\mathcal{X}}}$ for some $\overline{H}\in\widetilde{\mathrm{Pic}}(X/\mathbb{C})$ [Guo, Theorem 1.2].
@@
Remark 1.2.
We note that Conjecture 1.1 could be formulated verbatim for general families of polarized endomorphisms; see [GV1, §2.3] for a construction of the fibered Green current in this setting. We chose to keep the presentation concrete, since the seemingly more general statement reduces to the one given here by Fakhruddin’s observation in [Fa1], which shows that any polarized endomorphism extends to an endomorphism of a projective space.
Conjecture 1.1 is known to hold when $N=1$, because it follows from [De3, Theorem 1.1], as we explain in Section 3. In this article, we show that many existing works and conjectures can be viewed as special cases of Conjecture 1.1, sometimes in surprising ways. In fact, we will see that many powerful statements follow from a weaker form of Conjecture 1.1:
@@
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree >1>1, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible, flat family of subvarieties over $S$ containing a Zariski-dense set of preperiodic points of $\Phi$. Then $\mathcal{X}$ has codimension $\leq\dim S$ in a $\Phi$-special subvariety.
@@
Gao and Habegger pointed out in [GH2] that the converse assertion to Conjecture 1.3 is not true, with explicit examples and constructions in the setting where $\Phi$ fixes an abelian subscheme of $S\times\mathbb{P}^{N}$; see also [Gao]. This led them to formulate their [GH2, Theorem 1.3], which is what our Conjecture 1.1 aims to extend. Some of those counterexample constructions arose already in the (conjectural) characterization of subvarieties with geometric canonical height 0, in the Geometric Dynamical Bogomolov Conjecture formulated by Gauthier and Vigny [GV1]. The conjecture in [GV1] aims to extend theorems for abelian varieties over function fields of characteristic 0, proved in [GH1, CGHX]. We discuss the relation between [GV1, Conjecture 1.9] and Conjecture 1.1 in Section 5.
Remark 1.4.
In the case that both $\Phi$ and $\mathcal{X}$ are defined over $\overline{\mathbb{Q}}$, we expect Conjectures 1.1 and 1.3 to remain true upon replacing $\Phi$-preperiodic points by $\Phi$-small points, in the spirit of the Bogomolov Conjecture over number fields. Here a sequence of points $\{x_{n}\}_{n}\subset\mathcal{X}(\overline{\mathbb{Q}})$ is called $\Phi$-small if $\hat{h}_{\Phi}(x_{n})\to 0$, where $\hat{h}_{\Phi}(x):=\hat{h}_{\Phi_{\pi(x)}}(x)$ is the fiber-wise Call–Silverman canonical height over $\overline{\mathbb{Q}}$ introduced in [CS], and $\pi:S\times\mathbb{P}^{N}\to S$ is the projection. This would generalize the Relative Bogomolov Conjecture in [DGH3].
We conclude this article with a proof of the following implication.
@@
Contents of this article. We begin Section 2 by introducing a dynamical notion of $\Phi$-rank of $\mathcal{X}$, and we compare Conjecture 1.1 to the theorems of Gao and Habbeger from [GH2]. In particular, Conjecture 1.1 characterizes subvarieties $\mathcal{X}\subset S\times\mathbb{P}^{N}$ of maximal $\Phi$-rank; see Conjecture 2.3. We also compare the conditions of Conjecture 1.1 to the concept of non-degeneracy of subvarieties $\mathcal{X}$, which was originally introduced for families of abelian varieties. We then provide a brief survey in Section 3 of familiar cases of Conjecture 1.1, outside of the setting of abelian schemes. In particular, we remind the reader of the Dynamical Manin-Mumford Conjecture, first posed in [Zh], which corresponds to Conjecture 1.1 in the case where $\dim S=0$. We show that Conjecture 1.1 is known to hold in dimension $N=1$ over any base $S$ (and is in fact equivalent to [De3, Theorem 1.1]), and we relate it to the concept of $J$-stability for maps on $\mathbb{P}^{1}$. We then show that a conjecture of [BD2], and the recent classification of “special curves” in the moduli space $\mathrm{M}_{d}^{1}$ of maps on $\mathbb{P}^{1}$ in [JX], is a special case of Conjecture 1.1; see Theorem 3.5. Moreover, we illustrate the strength of Conjecture 1.1 with an example showing that uniform versions of the conjecture are consequences of the conjecture itself; the example we provide in Proposition 3.6 comes from the study of shared preperiodic points for distinct maps on $\mathbb{P}^{1}$. In Section 4, we show that Conjecture 1.1 implies the recent sparsity theorem of Gauthier, Taflin, and Vigny in [GTV] about PCF maps in the moduli spaces $\mathrm{M}_{d}^{N}$ of maps on $\mathbb{P}^{N}$, for $N>1$. We present Conjecture 4.1 as a special case of Conjecture 1.1 that extends the sparsity result of [GTV] to more general families of subvarieties in $\mathbb{P}^{N}$. We then explore the concept of minimal $\Phi$-rank and compare Conjecture 1.1 to the Geometric Bogomolov Conjecture posed in [GV1] in Section 5. Finally, in Section 6, we provide the proof of Theorem 1.5.
Acknowledgements. We would like to thank Ziyang Gao, Thomas Gauthier, Lars Kühne, Harry Schmidt, and Gabriel Vigny for many interesting discussions and their help during the preparation of this article. We thank the anonymous referees for helpful suggestions. We also thank the Simons Foundation for their support during a Symposium in August 2022 where this work was initiated. This project was supported in part with funding from the National Science Foundation, the Radcliffe Institute for Advanced Study, and the Natural Sciences and Engineering Research Council of Canada.
2. Dynamical rank and Betti rank
@@ 2.1. $\Phi$-Rank
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree >1>1, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible flat subvariety over $S$. We define the $\Phi$-rank of $\mathcal{X}$ to be
@@
For $\dim S=0$, so that $\Phi$ is a single endomorphism $f:\mathbb{P}^{N}\to\mathbb{P}^{N}$ defined over $\mathbb{C}$, and for any irreducible subvariety $Z$ of $\mathbb{P}^{N}$, we have
@@
$\int_{\mathbb{P}^{N}}\left(d^{-\ell n}(f^{n})^{*}\omega^{\wedge\ell}\right)\wedge[Z]\;=\;\int_{\mathbb{P}^{N}}\omega^{\wedge\ell}\wedge[Z]\;\not=\;0$
@@
We say that $\mathcal{X}\subset S\times\mathbb{P}^{N}$ has maximal $\Phi$-rank if $\operatorname{rank}_{\Phi}(\mathcal{X})=r_{\Phi,\mathcal{X}}$, the relative special dimension $r_{\Phi,\mathcal{X}}$ of $\mathcal{X}$, a name justified by the following proposition:
@@
$\dim_{S}\mathcal{X}\leq\operatorname{rank}_{\Phi}(\mathcal{X})\leq r_{\Phi,\mathcal{X}}.$
@@
Suppose that $\mathcal{Y}$ is $\Phi$-special, and set $k=\dim_{S}\mathcal{Y}$. Let $\Psi$ be a family of polarized endomorphisms over $S$ that commutes with an iterate of $\Phi$ along $\mathcal{Z}$ with $\mathcal{Y}\subset\mathcal{Z}\subset S\times\mathbb{P}^{N}$; replacing $\Psi$ with an iterate, we assume that $\mathcal{Y}$ is prefixed for $\Psi$. Now replace $\mathcal{Y}$ with an iterate under $\Psi$ so that $\Psi(\mathcal{Y})=\mathcal{Y}$, and pass to a normalization if it is not itself normal. We have $\hat{T}_{\Phi}=\hat{T}_{\Psi}$ on $\mathcal{Y}$. The current $\hat{T}_{\Psi}$, for a general family of polarized endomorphisms, is defined similarly to $\hat{T}_{\Phi}$; see [GV1, §2.3]. Since $(\Psi,\mathcal{Y})$ defines a family of polarized dynamical systems of degree $e>1$, there is a (1,1)-form $\omega$ on $\mathcal{Y}$ representing the first Chern class of the polarizing (relatively ample) line bundle $\mathcal{L}$ so that $\omega^{k+1}=0$. Because of the local-uniform convergence of the potentials of pullbacks of $\omega$ to a potential for $\hat{T}_{\Psi}$, we know that the wedge power $\hat{T}_{\Psi}^{\wedge j}$ is the limit of pullbacks $\frac{1}{e^{jn}}(\Psi^{n})^{*}(\omega^{\wedge j})$ on $\mathcal{Y}$, for any $j\geq 1$. As $\omega^{\wedge(k+1)}=0$, we have $\hat{T}_{\Phi}^{\wedge(1+\dim_{S}\mathcal{Y})}\wedge[\mathcal{Y}]=0$.
@@
For the lower bound on $\operatorname{rank}_{\Phi}(\mathcal{X})$, observe that the slice $(1,1)$-current $T_{s}$ of $\hat{T}_{\Phi}$ over $s\in S$ satisfies $T_{s}^{\wedge\ell}\wedge[X]\not=0$ for every subvariety $X$ of dimension $\ell$ in $\mathbb{P}^{N}$ and every $s\in S$, by Lemma 2.1. The nonvanishing of $\hat{T}_{\Phi}^{\wedge\dim_{S}\mathcal{X}}\wedge[\mathcal{X}]$ follows; see, for example, [BaBe, Proposition 4.3] on slicing. ∎
Condition (2) of Conjecture 1.1 means that $\operatorname{rank}_{\Phi}(\mathcal{X})\geq r_{\Phi,\mathcal{X}}$. By Proposition 2.2, Conjecture 1.1 is thus a characterization of subvarieties $\mathcal{X}$ of maximal $\Phi$-rank:
@@
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree >1>1, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible subvariety which is flat over $S$. The following are equivalent: